# Subgraph-Aware Training of Text-based Methods for Knowledge Graph Completion

### v3

@@ 5.3 Ablation study
BERT-base 0.433 0.389 0.476 0.560 220M
DeBERTa-base 0.434 0.388 0.478 0.559 280M
BERT-large 0.432 0.367 0.475 0.541 680M
DeBERTa-large 0.426 0.368 0.428 0.551 800M
Table 5: Performance comparison for different encoders.
5.4 Performance Across Encoders
To investigate the impact of the encoder architecture and the number of model parameters, we conduct experiments replacing BERT-base in SATKGC with BERT-large, DeBERTa-base, and DeBERTa-large (He et al., 2020). Table 5 presents the results. SATKGC is highly compatible with different encoders, showing the competitive performance. DeBERTa-large fine-tuned by SATKGC achieves the best performance on WN18RR. In addition, an increase in the number of model parameters may not necessarily result in enhanced performance on KGC, e.g., BERT-large on WN18RR, and BERT-large and DeBERTa-large on NELL-995 underperform the smaller encoders.
5.5 Comparing Sampling Methods
@@
RWR 0.682 0.620 0.728 0.821
BRWR 0.689 0.621 0.731 0.823
BRWR_P 0.673 0.610 0.726 0.810
MCMC 0.676 0.609 0.716 0.802
@@
RWR 0.355 0.264 0.382 0.532
BRWR 0.368 0.279 0.403 0.548
BRWR_P 0.347 0.258 0.376 0.526
MCMC 0.346 0.259 0.376 0.521
Table 6: Experiments comparing subgraph sampling methods on WN18RR and FB15k-237.
We investigate how model performance varies depending on the probability distribution $p_{v}$ used for neighbor selection in Section 4.1.888Recall that in our BRWR algorithm, a random walker selects one of the neighbors $v$ of a current node based on the inverse degree distribution $p_{v}$. We compare the performance of SATKGC using $p_{v}$ in subgraph sampling with two variants, one with $p_{v}$ replaced by the uniform distribution (dubbed RWR) and the other with $p_{v}$ replaced by the degree proportional distribution (dubbed BRWR_P). Table 6 shows the results. The three methods mostly outperform existing KGC models in Hits@1, with BRWR performing best. BRWR_P performs the worst, likely due to many duplicate entities in the extracted subgraphs sampled from the degree-proportional distribution. We also employ a Markov chain Monte Carlo (MCMC) based subgraph sampling method (Yang et al., 2020), referred to as MCMC (see details of MCMC in Appendix B). Note that in Table 2, MCMC outperforms other methods in terms of Hits@1 on WN18RR.
6 Analysis
6.1 Analysis on Negative Triples
Figure 5: Percentage of in-batch negative triples on FB15k-237 according to the range of cosine similarity scores predicted by SimKGC for different epochs and by SATKGC for different phases.
Figure 5 shows how the cosine similarity distribution of in-batch negative triples varies depending on the epoch of SimKGC, and the phase of SATKGC for FB15k-237. SATKGC encounters consistently more hard negatives with scores from 0.2 to 1.0 than SimKGC, though the majority of the scores range from -0.2 to 0.2 by the end of training for both methods. [^9] This trend is also observed in WN18RR. We speculate that SATKGC ends up with distinguishing positives from the hard negatives sampled from the subgraphs of a KG, as opposed to SimKGC which employs randomly-sampled easy negatives. The effectiveness using only negatives from SaaM in enhancing performance is demonstrated in Appendix G.
Based on our analysis, only 2.83% and 4.42% of the true triples ranked within the top 10 by SimKGC drop out of the top 10 for SATKGC on WN18RR and FB15k-237, respectively. In contrast, 34.87% and 13.03% of the true triples dropping out of the top 10 for SimKGC are ranked within the top 10 by SATKGC. This indicates that SATKGC effectively samples hard negatives while reducing false negatives.
6.2 Skewness of Occurrences
@@
Figure 6: (a) Number of occurrences of triples; (b) number of occurrences of entities. Both are counted throughout the entire training process for RANDOM and the SaaM variants, i.e., RWR, BRWR_P, and BRWR.
@@
Figures 6(a) and 6(b) show the number of visits of every triple and every entity, respectively, in the training process for RANDOM and the SaaM variants: RWR, BRWR_P, and BRWR. RANDOM, adopted in all the baselines, denotes selecting a mini-batch of triples at random without replacement. Triples and entities are sorted by descending frequency. Figure 6(a) demonstrates that RANDOM exhibits a uniform distribution, while RWR, BRWR_P, and BRWR display varying degrees of skewness, with BRWR being the most skewed and BRWR_P the least. Figure 6(b) illustrates that BRWR is the least skewed whereas RANDOM shows the most skewed distribution. As a result, a larger skewness in the distribution of number of visits for triples in turn leads to more equally visiting entities, thus improving the performance. [^10] The similar trends are shown on WN18RR.
Further analysis reinforces this finding. In FB15k-237, the average degree of FP triples’ tails is 75 for SATKGC and 63 for SimKGC. A smaller portion of low-degree tails for SATKGC than for SimKGC indicates that exposure to more low-degree entities $\hat{t}$ in training helps the model position their embeddings farther from the $(h,r)$ embeddings for negative triples $(h,r,\hat{t})$, as SATKGC visits low-degree entities more often during training than RANDOM for SimKGC. [^11] This is also confirmed in all the other datasets.
@@
For BRWR in Figure 6(a), we examine the structural characteristics on sets $S_{m}$ and $S_{l}$ of entities in $1,000$ most and least frequent triples, respectively. The entities in $S_{m}$ have an average degree of $11.1$, compared to $297.3$ for those in $S_{l}$. The betweenness centrality [^12] The betweenness centrality of node $v$ is the number of shortest paths that pass through $v$ in the graph divided by the total number of shortest paths between all pairs of nodes, which measures $v$’s importance in the graph. averages around $5.2\times 10^{-5}$ for $S_{m}$ and $8.2\times 10^{-4}$ for $S_{l}$. These observations implies that SaaM prioritizes visiting unpopular and peripheral triples in the KG over focusing on information-rich triples.
7 Conclusion
@@
In this paper, we propose a generic training scheme and a new contrastive learning method for KGC. Harmonizing (i) SaaM using a subgraph of KG as a mini-batch, and (ii) contrastive learning that incorporates the structural hardness of the KG into fine-tuning PLMs helps learning the contextual text embeddings aware of the difficulty in the structural context of the KG. Our findings imply that unequally feeding triples in training and leveraging the unique characteristics of KG lead to the effective text-based KGC method, achieving state-of-the-art performance on the four KG benchmarks.
Limitations
While our training scheme efficiently samples many hard negatives from subgraphs, BRWR incurs a computational overhead to extract subgraphs from Wikidata5M before training (see Appendix E). However, note that this algorithm can be efficiently parallelized or replaced by a simpler one.
@@ Appendix A Inference
For inference, the bi-encoder model calculates the cosine similarity between $\textbf{x}_{hr}$ for a given $(h,r,?)$ and $\textbf{x}_{t}$ for all entities $t$. Then the tails with the top-k largest cosine similarities are answers. For a single pair, we need $|E|$ forward passes of $\textit{Encoder}_{t}$ to obtain $\textbf{x}_{t}$ for all entities $t$. Given a set $T$ of test triples, $2|T|$ forward passes of $\textit{Encoder}_{hr}$ are required to get $\textbf{x}_{hr}$ for every triple $(h,r,?)\in T$ and its inverse triple $(t,r^{-1},?)$, thus resulting in $O(|E|+|T|)$ computation in total.
@@ Appendix B Markov Chain Monte Carlo Based Subgraph Sampling
Inspired by a negative sampling approach (Yang et al., 2020) based on Markov chain Monte Carlo (MCMC) in a general graph, we propose a new method to sample subgraphs from a KG. A negative sampling distribution should be positively but sublinearly correlated with the positive sampling distribution, which was validated by Yang et al. (2020). To include the entities close to a positive triple in the KG in negative triples, we define the sampling distribution $p_{n}$ of the negative tail $\hat{t}$ as : $p_{n}(\hat{t}|h,r)\propto p_{d}(\hat{t}|h,r)^{\alpha},0<\alpha<1$, $p_{d}(\hat{t}|h,r)=\frac{\cos(\textbf{x}_{hr},\textbf{x}_{\hat{t}})}{\sum_{e\in E}\cos(\textbf{x}_{hr},\textbf{x}_{e})}$. where $\alpha$ is a parameter to stabilize the optimization process, $E$ is a set of entities in the KG, and $p_{d}$ is the sampling distribution of the positive tail. Calculating the normalization term in $p_{d}(\hat{t}|h,r)$ is time consuming and almost impossible. Therefore, we sample a negative tail $\hat{t}$ from $\tilde{p_{d}}=\cos(\textbf{x}_{hr},\textbf{x}_{\hat{t}})$ by using the Metropolis-Hastings (M-H) algorithm, and randomly select a triple whose head is $\hat{t}$.
@@ Appendix C Datasets
In this paper, we use four KGC benchmarks. WN18RR is a sparse KG with a total of 11 relations and $\sim 41k$ entities. WN18RR is the dataset derived from WN18, consisting of relations and entities from WordNet Miller (1992). WN18RR addresses the drawbacks of test set leakage by removing the inverse relation in WN18. NELL-995 is a sparse KG extracted from the web. FB15k-237 is a dense KG with 237 relations. Wikidata5M, a much larger KG than the others, provides transductive and inductive settings. Wikidata5M-Trans is for the transductive setting, where entities are shared and triples are disjoint across training, validation, and test. Wikidata5M-Ind is for the inductive setting, where the entities and triples are mutually disjoint across training, validation, and test (Wang et al., 2021c).
Appendix D Implementation Details
In our weighted InfoNCE loss, additive margin $\gamma$ is set to 0.02. We select the best performing batch sizes of 1024 from 512, 1024, 1536, 2048 for WN18RR, NELL-995, and Wikidata5M, and 3072 from 1024, 2048, 3072, 4096 for FB15k-237. We set the restart probability $p_{r}$ to 1/25 in BRWR. We used six A6000 GPUs and 256G RAM. Training on WN18RR took 50 phases, for a total of 4 hours. NELL-995 and FB15k-237 took 20 phases and a total of 3 hours and 10 hours respectively, while Wikidata5M took 2 phases and 20 hours.
@@ Appendix D Implementation Details
1024 0.322 0.234 0.347 0.500
2048 0.339 0.250 0.368 0.521
3072 0.368 0.279 0.400 0.548
4096 0.357 0.261 0.357 0.532
Table 7: The results of investigating the model performance with respect to the batch size.
Appendix E Runtime Analysis
Sampling Time
Algorithm WN18RR FB15k-237 Wiki5M-Trans Wiki5M-Ind
BRWR 13m 1h 14h 14h
Shortest Path + Degree 7m 15m 7h 7h
Training Time (SATKGC: time per phase, others: time per epoch)
Model WN18RR FB15k-237 Wiki5M-Trans Wiki5M-Ind
SATKGC 7m 30m 10h 10h
SimKGC Wang et al. (2022a) 4m 20m 9h 9h
StAR Wang et al. (2021a) 1h 1h 30m - -
SANS Ahrabian et al. (2020) 45m 57m - -
Table 8: The elapsed time required for sampling and the time allotted per phase during training.
Training SATKGC incurs a marginal computational overhead because (i) sampling subgraphs and (ii) computing distances and degrees are performed in advance before training. As shown in Table 8, the computational cost for (i) and (ii) is acceptable. The cost depends on the size of a mini-batch, which can be adjusted. Moreover, the time complexity for (ii) is acceptable. For each mini-batch $B$ of triples, we run Dijkstra’s single source shortest path algorithm, and thus the runtime to compute the distance is O($|V|log|V|+|B|log|V|$), where $V$ is a set of entities in $B$. As described in Section 4.3.1, we do not calculate all-pair shortest paths for every pair of vertices in $V$. To reduce the computational overhead, we compute the approximate distance by using the shortest path between the head of the center triple in a subgraph and every tail in that subgraph, which have been already obtained fro Dijkstra’s algorithm. Table 8 shows the training time per phase for SATKGC and per epoch for SimKGC Wang et al. (2022a), StAR Wang et al. (2021a), and SANS Ahrabian et al. (2020). SATKGC remains competitive, though it takes slightly more time than SimKGC due to the computational overhead for (a) computing its loss using the shortest path weight (SPW) and degree weight (DW), (b) counting the occurrences of visited triples to select the next subgraph, and (c) fetching subgraphs. StAR and SANS take longer than SATKGC. StAR runs out of memory on the Wikidata5M datasets, while SANS cannot be applied to the inductive setting (Wikidata5M-Ind), and is not expected to finish within a reasonable time on Wikidata5M-Trans.

### v5

@@ 5.3. Ablation Study
MPNet 0.370 0.279 0.403 0.550 220M
BERT-base 0.366 0.273 0.400 0.546 220M
DeBERTa-base 0.359 0.274 0.399 0.545 280M
BERT-large 0.339 0.268 0.378 0.532 680M
DeBERTa-large 0.345 0.272 0.389 0.540 800M
Table 5. Performance comparison for different encoders.
5.4. Performance Across Encoders
To investigate the impact of the encoder architecture and the number of model parameters, we conduct experiments replacing MPNet in SATKGC with BERT-base, BERT-large, DeBERTa-base, and DeBERTa-large (He et al., 2020). Table 5 presents the results. SATKGC is highly compatible with different encoders, showing the competitive performance. DeBERTa-large fine-tuned by SATKGC achieves the best performance on WN18RR. In addition, an increase in the number of model parameters may not necessarily result in enhanced performance on KGC111111e.g., BERT-large on WN18RR, and BERT-large and DeBERTa-large on FB15k-237 underperform the smaller encoders..
5.5. Comparing Subgraph Sampling Methods
@@
RWR 0.690 0.622 0.730 0.825
BRWR 0.694 0.626 0.743 0.833
BRWR_P 0.676 0.615 0.727 0.823
MCMC 0.687 0.609 0.736 0.825
@@
RWR 0.358 0.269 0.385 0.534
BRWR 0.370 0.279 0.403 0.550
BRWR_P 0.351 0.262 0.383 0.530
MCMC 0.332 0.245 0.362 0.507
Table 6. Experiments comparing subgraph sampling methods on WN18RR and FB15k-237.
@@
Figure 6. (a) Number of occurrences of triples; (b) number of occurrences of entities. Both are counted throughout the entire training process for RANDOM and the SaaM variants, i.e., RWR, BRWR_P, and BRWR.
We investigate how model performance varies depending on the probability distribution $p_{v}$ used for neighbor selection in Section 4.1.Recall that in our BRWR algorithm, a random walker selects one of the neighbors $v$ of a current node based on the inverse degree distribution $p_{v}$. We compare the performance of SATKGC using $p_{v}$ in subgraph sampling with two variants, one with $p_{v}$ replaced by the uniform distribution (dubbed RWR) and the other with $p_{v}$ replaced by the degree proportional distribution (dubbed BRWR_P). Table 6 shows the results. The three methods mostly outperform existing KGC methods in Hits@1, with BRWR performing best and BRWR_P performing worst. We also employ a Markov chain Monte Carlo (MCMC) based subgraph sampling method (Yang et al., 2020), referred to as MCMC (see details of MCMC in Appendix B). Note that in Table 2, MCMC outperforms other methods for Hits@1 on WN18RR.
@@
To verify why variations in $p_{v}$ lead these differences, we conduct further analysis. Figures 6(a) and 6(b) show the number of visits of every triple and every entity, respectively, in the training process for RANDOM and the SaaM variants, i.e., RWR, BRWR_P, and BRWR. RANDOM, adopted in all the baselines denotes selecting a mini-batch of triples at random without replacement. Triples and entities are sorted by descending frequency. Figure 6(a) demonstrates that RANDOM exhibits a uniform distribution, while RWR, BRWR_P, and BRWR display varying degrees of skewness, with BRWR being the most skewed and BRWR_P the least. Figure 6(b) illustrates that BRWR is the least skewed whereas RANDOM shows the most skewed distribution. BRWR_P, which employs the degree-proportional distribution, extracts many duplicate entities in the subgraphs, leading to the most skewed distribution among the SaaM variants. As a result, a larger skewness in the distribution of number of visits for triples in turn leads to more equally visiting entities, thus improving the performance. [^12] The similar trends are shown on WN18RR. Further analysis reinforces this finding. In FB15k-237, the average degree of FP triples’ tails is 75 for SATKGC and 63 for SimKGC. A smaller portion of low-degree tails for SATKGC than for SimKGC indicates that exposure to more low-degree entities $\hat{t}$ in training helps the model position their embeddings farther from the $(h,r)$ embeddings for negative triples $(h,r,\hat{t})$, as SATKGC visits low-degree entities more often during training than RANDOM for SimKGC. [^13] This is also confirmed in all the other datasets.
@@
We examine the structural characteristics on sets $S_{m}$ and $S_{l}$ of entities in $1,000$ most and least frequent triples, respectively, visited by BRWR. The entities in $S_{m}$ have an average degree of $11.1$, compared to $297.3$ for those in $S_{l}$. The betweenness centrality, which quantifies a node’s importance by measuring the fraction of shortest paths passing through it, averages around $5.2\times 10^{-5}$ for entities in $S_{m}$ and $8.2\times 10^{-4}$ for $S_{l}$.These observations implies that SaaM prioritizes visiting unpopular and peripheral triples in KG over focusing on information-rich triples.
Figure 7. Percentage of in-batch negative triples on FB15k-237 according to the range of cosine similarity scores predicted by SimKGC and SATKGC for different epochs.
5.6. Analysis on Negative Triples
Figure 7 shows how the cosine similarity distribution of in-batch negative triples varies depending on the epoch of SimKGC and SATKGC for FB15k-237. SATKGC encounters consistently more hard negatives with scores from 0.2 to 1.0 than SimKGC, though the majority of the scores range from -0.2 to 0.2 by the end of training for both methods. [^14] This trend is also observed in WN18RR. We speculate that SATKGC ends up with distinguishing positives from the hard negatives sampled from the subgraphs of the KG, as opposed to SimKGC which employs randomly-sampled easy negatives.
Based on our analysis, only 2.83% and 4.42% of the true triples ranked within the top 10 by SimKGC drop out of the top 10 for SATKGC on WN18RR and FB15k-237, respectively. In contrast, 34.87% and 13.03% of the true triples dropping out of the top 10 for SimKGC are ranked within the top 10 by SATKGC. This indicates that SATKGC effectively samples hard negatives while reducing false negatives. The additional experiment on in-batch negatives is provided in Appendix H.
6. CONCLUSION
@@
We propose a generic training scheme and new contrastive learning (CL) for KGC. Harmonizing SaaM using a subgraph of KG as a mini-batch, and both CL and mini-batch training incorporating the structural inductive bias of KG in fine-tuning PLMs helps learning the contextual text embeddings aware of the difficulty in the structural context of KG. Our findings imply that unequally feeding triples in training and leveraging the unique characteristics of KG lead to the effective text-based KGC method, achieving state-of-the-art performance on the three KG benchmarks.
Acknowledgements.
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2020-II201373, Artificial Intelligence Graduate School Program(Hanyang University)).
@@ Appendix A Inference
For inference, we calculate the cosine similarity between $\textbf{x}_{hr}$ for a given $(h,r,?)$ and $\textbf{x}_{t}$ for all entities $t$. Then the tails with the top-k largest cosine similarities are answers. For a single pair, we need $|E|$ forward passes of $\textsf{Encoder}_{t}$ to obtain $\textbf{x}_{t}$ for all entities $t$. Given a set $T$ of test triples, $2|T|$ forward passes of $\textsf{Encoder}_{hr}$ are required to get $\textbf{x}_{hr}$ for every triple $(h,r,?)\in T$ and its inverse triple $(t,r^{-1},?)$, thus resulting in $O(|E|+|T|)$ computation in total.
@@ Appendix B Markov Chain Monte Carlo Based Subgraph Sampling
Inspired by a negative sampling approach (Yang et al., 2020) based on Markov chain Monte Carlo (MCMC) in a general graph, we propose a new method to sample subgraphs from the KG. A negative sampling distribution should be positively but sublinearly correlated with the positive sampling distribution, which was validated by Yang et al. (2020). To include the entities close to a positive triple in the KG in negative triples, we define the sampling distribution $p_{n}$ of the negative tail $\hat{t}$ as : $p_{n}(\hat{t}|h,r)\propto p_{d}(\hat{t}|h,r)^{\alpha},0<\alpha<1$, $p_{d}(\hat{t}|h,r)=\frac{\cos(\textbf{x}_{hr},\textbf{x}_{\hat{t}})}{\sum_{e\in E}\cos(\textbf{x}_{hr},\textbf{x}_{e})}$. where $\alpha$ is a parameter to stabilize the optimization process, $E$ is a set of entities in the KG, and $p_{d}$ is the sampling distribution of the positive tail. Calculating the normalization term in $p_{d}(\hat{t}|h,r)$ is time consuming and almost impossible. Therefore, we sample a negative tail $\hat{t}$ from $\tilde{p_{d}}=\cos(\textbf{x}_{hr},\textbf{x}_{\hat{t}})$ by using the Metropolis-Hastings (M-H) algorithm, and randomly select a triple whose head is $\hat{t}$.
@@ Appendix C Datasets
In this paper, we use three KGC benchmarks. WN18RR is a sparse KG with a total of 11 relations and $\sim 41k$ entities. WN18RR is the dataset derived from WN18, consisting of relations and entities from WordNet (Miller, 1995). WN18RR addresses the drawbacks of test set leakage by removing the inverse relation in WN18. FB15k-237 is a dense KG with 237 relations. Wikidata5M, a much larger KG than the others, provides transductive and inductive settings. Wikidata5M-Trans is for the transductive setting, where entities are shared and triples are disjoint across training, validation, and test. Wikidata5M-Ind is for the inductive setting, where the entities and triples are mutually disjoint across training, validation, and test (Wang et al., 2021a).
Appendix D Input Format Details for Encoders
We adopt MPNet (Song et al., [n. d.]) as Encoderhr and Encodert, and set the maximum token length to 50. For each triple $h,r,t$, we concatenate the name of $h$, its textual description, and the name of $r$ with a special symbol [SEP] in between, and treat the concatenation as input for Encoderhr, while the input for Encodert consists the name and textual description of $t$. As illustrated in Figure 8, given a triple (Leonardo da Vinci, painted, Mona Lisa), the input to Encoderhr is the concatenation of [CLS], ”Leonardo da Vinci”, [SEP], the head’s description, [SEP], and the relation’s name ”painted.” The input to Encodert is the concatenation of [CLS], ”Mona Lisa”, [SEP], and the tail description.
Appendix E Implementation Details
In our weighted InfoNCE loss, additive margin $\gamma$ is set to 0.02. We select the best performing batch sizes of 1024 from {512, 1024, 1536, 2048} for WN18RR, and Wikidata5M, and 3072 from {1024, 2048, 3072, 4096} for FB15k-237. We set the restart probability $p_{r}$ to 1/25 in BRWR. We used six A6000 GPUs and 256G RAM. Training on WN18RR took 50 epochs, for a total of 4 hours. FB15k-237 took 30 epochs and a total of 3 hours, while Wikidata5M took 2 epochs and 19 hours.
@@ Appendix E Implementation Details
1024 0.337 0.250 0.366 0.519
2048 0.352 0.255 0.374 0.525
3072 0.370 0.279 0.403 0.550
4096 0.362 0.270 0.391 0.542
Table 7. The results of investigating the model performance with respect to the batch size.
Figure 9. (a) Impact of varying restart probabilities on the model performance. (b) Impact of varying the number of duplicate entities in a mini-batch on the model performance.
Appendix F Hyperparameter Sensitivity
We investigate how restart probability $p_{r}$ in BRWR affects model performance. The hyperparameter $p_{r}$ is associated with the length of the random walk path from the start entity, which in turn influences the occurrence of duplicate entities in a mini-batch. A longer path leads to fewer duplicate entities in the mini-batch. Figure 9(a) illustrates that a lower $p_{r}$ value, encouraging a longer random walk path, leads to higher Hits@1 for WN18RR and FB15k-237. We analyze the impact of duplicate entities in a mini-batch on the model performance. In Figure 9(b), more duplicate entities resulting from higher $p_{r}$ negatively impact on the performance, which highlights the importance of reducing the duplicates in a mini-batch to avoid the performance degradation. Additionally, we compare the performance of SATKGC on WN18RR and FB15k-237 with varying batch sizes in Table 7. Based on the results, we select the batch size that produce the best performance as the default setting for SATKGC.
Appendix G Performance of the SaaM Scheme Across Models
To demonstrate the generality of the SaaM scheme, we conduct additional experiments by applying the SaaM approach to RotatE, an embedding-based method, and StAR, a text-based method, to evaluate their performance on the WN18RR and FB15k-237 datasets. As shown in Table 8, incorporating the SaaM approach into both RotatE and StAR results in significant improvements across all metrics. For instance, on the FB15k-237, StAR + SaaM exhibits a significant improvement in Hits@1, increasing by 12.8%, from 0.195 to 0.220. These findings illustrate that SaaM is model-agnostic and highly generalizable, effectively enhancing performance regardless of the underlying model.